OpenStudy (anonymous):

During the annual shuffleboard competition, Renee gives her puck an initial speed of 9.32 m/s. Once leaving her stick, the puck slows down at a rate of -4.06 m/s/s. a. Determine the time it takes the puck to slow to a stop. b. Use your initial speed and the calculated time to determine the average speed and the distance which the puck travels before stopping.

OpenStudy (anonymous):

a. Use the cinematic equation V = Vo + at. You have Vo (9.32m/s), you have a (-4.06m/s^2) and V = 0 (stop). Solve for t. b. Try to work it out now. :-) If you still can't, ask again

OpenStudy (unklerhaukus):

*kinematic

OpenStudy (anonymous):

@UnkleRhaukus Indeed, I didn't intend to talk about movies :-) My bad.

OpenStudy (anonymous):

i got the first one a) 2.30 s, but having trouble with part b

OpenStudy (anonymous):

a) a=-4.06 m/s^2. u=9.32 m/s v=0 m/s t=? a=(v-u)/t or, t=(v-u)/a thus, t= (9.32/4.06) = 2.3 sec. b) average speed = (u + v)/2 thus, avg speed = 4.66 m/s And, distance traversed before stopping: using eq: s=((u+v)/2)*t s= 10.72m..

OpenStudy (anonymous):

ok nice work thanx

OpenStudy (anonymous):

welcome...:)